OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. n^4-11n^2+30/n^4-7n^2+10

OpenStudy (anonymous):

\[\frac{ n^{4}-11n^{2}+30 }{n^{4}-7n ^{2}+10 }\]

OpenStudy (anonymous):

did you try to factor?

OpenStudy (anonymous):

idk how to

OpenStudy (anonymous):

Are you familiar with AC method of factoring?

OpenStudy (anonymous):

nope :/

OpenStudy (anonymous):

AC method is mostly used for factoring quadratic expressions. But in this special case, both the numerator and denominator are very similar to quadratic expressions. So the method is can be applied. Just pretend x^2 = u, then you have, u^2 - 11u + 30, do you see what I did there?

OpenStudy (anonymous):

ohh. yes i see

OpenStudy (anonymous):

ok, now it's just the matter of factoring u^2 - 11u + 30 using AC method. Recall that the general quadratic equations is Ax^2 + Bx + C. But for sake of matching variables, I'm going to change it to Au^2 + Bu + C Ok so far?

OpenStudy (anonymous):

ive got this so far...i did n^2=t so..

OpenStudy (anonymous):

\[\frac{ t^{2}-11t+30 }{ t^{2}-7t+10 }\]

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

\[\frac{ (t-5)(t-6) }{ (t-2)(t-5) }\]

OpenStudy (anonymous):

(^-^)b good job. Save me time haha. Now do you see anything canceling?

OpenStudy (anonymous):

\[\frac{ n^{2} -6}{ n^{2}-2}\]

OpenStudy (anonymous):

now i just need to figure out the restrictions

OpenStudy (anonymous):

well, can you divide by 0?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

right. What does that tell you about the denominator?

OpenStudy (anonymous):

idk >_< im lost again

OpenStudy (anonymous):

can denominator be 0?